The percentage by volume of C3H8 in a gaseous mixture of C3H8,CH4 and Co2 is 20.Calculate the volume of Co2 produced by complete combustion under similar condition
Ans-140ml

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Kaushaki
 
Give the blend a chance to contain C3H8,CH4 ,CO as a,b,c . 
a+b+c=100ml 
b+c=100-20 
at the point when consumed in abundance of air, 
C3H8+5O2=3CO2+5H2O 
CH4+2CO2=CO2+2H2O 
CO+1/2O2=CO2 
CO2 framed regarding mole=3a+b+c 
=3*20+(100-20) 
=60+80 
=140ml